OpenStudy (anonymous):

Choose the correct simplification of (x^4/y^2)^4

OpenStudy (theviper):

\[\large{(\dfrac{x^4}{y^2})^4}\]\[\large{=\dfrac{(x^4)^4}{(y^2)^4}}\]\[=\Large{\dfrac{x^{4\times4}}{y^{2\times4}}}\]\[\large{=??}\]

OpenStudy (hba):

See this example \[\huge\ \ ( \frac{ a^2 }{ b^2 })^2=\frac{ a^{2 \times 2} }{ b^{2 \times 2} }\]

OpenStudy (anonymous):

x^8/y^6?

OpenStudy (hba):

No it would be \[x^{4 \times 4}/y^{2\times 4}\] Don't add,multiply.

OpenStudy (theviper):

yep :)

OpenStudy (anonymous):

x^16/y^8?

OpenStudy (hba):

Yeah you nailed it right :)

OpenStudy (theviper):

yes

OpenStudy (anonymous):

you both give awsome responses, is there a way to add you botincase i might need help in the future?

OpenStudy (hba):

Yes you can use @ and then the name like @hba